# On Fock covariance for product systems and the reduced Hao–Ng isomorphism problem by discrete actions

### v1

@@ 2.4. Product systems
Associativity of the product gives that $M_{\xi_{p}}^{q,pq}\in{\mathcal{L}}(X_{q},X_{pq})$ with
@@
where the sum is taken in the s*-topology, and with the understanding that we are embedding ${\mathcal{L}}(X_{q},X_{pq})\hookrightarrow{\mathcal{L}}({\mathcal{F}}X)$ as the $(pq,q)$-entry. By applying on $\eta_{r}\in X_{r}\subseteq{\mathcal{F}}X$ we derive that
@@
We start with the following three comments. First, suppose that $s=p^{-1}r$ for some $p,s\in P$. Then by definition we have that
@@
Next, suppose that $s=qr$ for some $q,s\in P$. Then by definition we have that
@@
Finally, suppose that $s=p^{-1}qr$ for some $p,q,s\in P$. Then we have that
@@
We proceed with the proof. Let $r$ be in $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ for some $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. We have that there is an $s_{1}\in q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{1}=p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r$. From the comments above we then have that $\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\in X_{s_{1}}$ and
@@
Continuing inductively for each $k=3,\dots,n-1$ we obtain an $s_{k}\in q_{n-k}^{-1}p_{n-k}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{k}=p_{n-(k-1)}^{-1}q_{n-(k-1)}s_{k-1}$, and for each $k=1,\dots,n-1$ we have that
@@
For the Fock representation we also have the following proposition.
@@
If $r\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $a\in A:=X_{e}$, then by Proposition 2.3 we have that
@@
for $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ and $n\in\mathbb{N}$. A proof can be found in [47, Lemma 2.2] for non-degenerate product systems, but similar arguments give the conclusion in the general case.
@@
We write $\mathbf{K}_{{\mathbf{x}},t_{\ast}}$ for this closed space. We do not claim that $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$. Note that, if $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$, then
@@
Indeed, since $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ we have that
@@
with $r_{1}^{-\varepsilon_{2}}s_{1}\cdots r_{m}^{-1}s_{m}^{\varepsilon_{2}^{\prime}}p_{1}^{-\varepsilon_{1}}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon_{1}^{\prime}}=e$. Hence we obtain that
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $t$ be a representation of $X$ and let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. For $\emptyset\neq{\mathbf{x}}\in{\mathcal{J}}$ and $r\in{\mathbf{x}}$ we have that
@@
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $r\in{\mathbf{x}}$. By Proposition 2.3 and Proposition 2.4 we have that
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. Then for every ${\mathbf{x}}\in{\mathcal{J}}$ and $r\not\in{\mathbf{x}}$ we have that
@@
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $\xi_{r}\in X_{r}$. By Proposition 2.4 we obtain that
@@
By the universal property, the Toeplitz C*-algebra ${\mathcal{T}}(X)$ admits a coaction by $G$. We let
@@
On the other hand, by [17, Theorem 19.5] we obtain a *-epimorphism
@@
As noted in [11, Proposition 4.1], the representation $\lambda$ admits a reduced coaction by using the unitary
@@
Indeed, for this $U$ we have that
@@ 2.5. Fowler’s product systems
Fowler [18] defines a product system $X$ over $P$ with coefficients in a C*-algebra $A$ as a family $\{X_{p}\}_{p\in P}$ of C*-correspondences over $A$ together with a family of unitary maps $\{u_{p,q}\colon X_{p}\otimes_{A}X_{q}\to X_{pq}\}_{p,q\in P}$ such that:
@@
Note here that item (ii) implies in particular that every $X_{p}$ is non-degenerate. To allow for comparisons, we will refer to $(X,\{u_{p,q}\}_{p,q\in P})$ as a product system in the sense of Fowler.
It can be directly verified that, if $X$ is a concrete product system in some ${\mathcal{B}}(H)$ such that $[X_{p}X_{q}]=X_{pq}$ for all $p,q\in P$, then the family $\{X_{p}\}_{p\in P}$ defines a product system in the sense of Fowler by considering the unitary maps $\{u_{p,q}\}_{p,q\in P}$ given by
@@
In this case the representations of $X$ are in bijection with the representations of $Y$. The reader may refer to [9] for the full details.
@@ 3.1. Fock covariant representations
defined by the coaction by $G$ on ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$. A representation of $X$ that promotes to a representation of ${\mathcal{F}}{\mathcal{C}}X$ will be called a Fock covariant representation of $X$. We note that, the definition of the Fock covariant bundle does not depend on the choice of $G$. This follows by applying the first part of the proof of [47, Lemma 3.9] in this setting. Note also that since the bimodule properties are graded, every representation of ${\mathcal{F}}{\mathcal{C}}X$ is a representation of $X$.
@@
By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we have that
@@
If $t$ is Fock covariant, then by definition $t_{\ast}$ factors through the canonical *-epimorphism $q_{\operatorname{c}}^{\operatorname{F}}\colon{\mathcal{T}}(X)\to{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$, and thus $\ker q_{\operatorname{c}}^{\operatorname{F}}\subseteq\ker t_{\ast}$. By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we also have that $\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}$, and therefore
@@
For the converse, recall that the fixed point algebra $[{\mathcal{T}}(X)]_{e}$ is the inductive limit of $\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for $\cap$-closed families ${\mathcal{F}}\subseteq{\mathcal{J}}$. By the properties of the inductive limits we have that
@@
Recall that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ by Proposition 2.6. Therefore, if $r\notin\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then we have that
@@
where in the first equality we used that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. By injectivity of $\lambda_{r}$, we have that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. This concludes that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$, as required.
Now suppose that $t$ is Fock covariant. Recall that
@@
On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ then for every $\xi_{r}\in X_{r}$ we have that
@@
where in the first equality we used that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. Injectivity of $t$ implies that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. We conclude that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$ and thus
@@
For the converse, suppose that $t$ satisfies conditions (i) and (ii). By Proposition 3.1, it suffices to prove that $\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\subseteq\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for every finite $\cap$-closed ${\mathcal{F}}\subseteq{\mathcal{J}}$. If ${}{\mathcal{F}}=\{\emptyset\}$, then condition (i) for $\lambda$ and $t$ yields
@@
where we used that $\mathbf{K}_{\emptyset,\hat{t}_{*}}=\mathbf{B}_{\{\emptyset\},\hat{t}_{\ast}}$ is a C*-subalgebra of $[{\mathcal{T}}(X)]_{e}$. Next suppose that =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ such that $\bigcup_{i=1}^{n}{\mathbf{x}}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$ such that
@@
For $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have that
@@
where in the first equality we used Proposition 2.5, and on the second equality we used that $\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$ if $r\notin\mathbf{x}_{i}$ from Proposition 2.6. Condition (iii) now yields that $\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$, and thus

### v2

@@ 2.4. Product systems
In his seminal paper, Fowler [20] formalised product systems motivated by the work of Nica [46], Pimsner [49], as well as his work with Raeburn [21]. However this definition imposes further conditions on the family of C*-correspondences. In this subsection we consider a more general framework that still allows the structural constructions relevant to product systems to be formulated yet without these extra conditions. In Subsection 2.5 we will give the connection with, and the comparison to, the product systems in the sense of Fowler.
@@
Associativity of the product yields $M_{\xi_{p}}^{q,pq}\in{\mathcal{L}}(X_{q},X_{pq})$ with
@@
where the sum is taken in the s*-topology, and with the understanding that we are embedding ${\mathcal{L}}(X_{r},X_{pr})\hookrightarrow{\mathcal{L}}({\mathcal{F}}X)$ as the $(pr,r)$-entry. By applying on $\eta_{r}\in X_{r}\subseteq{\mathcal{F}}X$ we derive that
@@
We start with the following three comments. First, suppose that $s=p^{-1}r$ for some $p,s\in P$. Then by definition we have
@@
Next, suppose that $s=qr$ for some $q,s\in P$. Then by definition we have
@@
Finally, suppose that $s=p^{-1}qr$ for some $p,q,s\in P$. Then we have
@@
We proceed with the proof. Let $r$ be in $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ for some $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. We have that there is an $s_{1}\in q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{1}=p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r$. From the comments above we then have $\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\in X_{s_{1}}$ and
@@
Continuing inductively for each $k=3,\dots,n-1$ we obtain an $s_{k}\in q_{n-k}^{-1}p_{n-k}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{k}=p_{n-(k-1)}^{-1}q_{n-(k-1)}s_{k-1}$, and for each $k=1,\dots,n-1$ we have
@@
For the Fock representation we have the following proposition, see also [20, page 340].
@@
If $r\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $a\in A:=X_{e}$, then by Proposition 2.3 we have
@@
for $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ and $n\in\mathbb{N}$. A proof can be found in [52, Lemma 2.2] for non-degenerate product systems, but similar arguments give the conclusion in the general case.
@@
We write $\mathbf{K}_{{\mathbf{x}},t_{\ast}}$ for this closed linear space. We do not claim that $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$. Note that, if $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$, then
@@
Indeed, since $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ we have
@@
with $r_{1}^{-\varepsilon_{2}}s_{1}\cdots r_{m}^{-1}s_{m}^{\varepsilon_{2}^{\prime}}p_{1}^{-\varepsilon_{1}}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon_{1}^{\prime}}=e$. Hence we obtain
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $t$ be a representation of $X$ and let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. For $\emptyset\neq{\mathbf{x}}\in{\mathcal{J}}$ and $r\in{\mathbf{x}}$ we have
@@
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $r\in{\mathbf{x}}$. By Proposition 2.3 and Proposition 2.4 we have
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. Then for every ${\mathbf{x}}\in{\mathcal{J}}$ and $r\not\in{\mathbf{x}}$ we have
@@
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $\xi_{r}\in X_{r}$. By Proposition 2.4 we obtain
@@
By the universal property, the Toeplitz C*-algebra ${\mathcal{T}}(X)$ admits a coaction by $G$, see for example [52, Lemma 2.2]. We let
@@
On the other hand, by [19, Theorem 19.5] we obtain a *-epimorphism
@@
As noted in [13, Proposition 4.1], the representation $\lambda$ admits a reduced coaction by using the unitary
@@
Indeed, for this $U$ we have
@@ 2.5. Fowler’s product systems
In [20, Definition 2.1] Fowler defines a product system $X$ over $P$ with coefficients in a C*-algebra $A$ as a family $\{X_{p}\}_{p\in P}$ of C*-correspondences over $A$ together with a family of maps $\{u_{p,q}\colon X_{p}\otimes_{A}X_{q}\to X_{pq}\}_{p,q\in P}$ such that:
@@
Brownlowe, Larsen and Stammeier note in [7, Remark 6.2] that if the group of units of the semigroup is non-trivial, then every $X_{p}$ is automatically non-degenerate, see also [37, Remark 1.3]. Hence it has been accustomed to work within the non-degenerate setup. To allow for comparisons, we will refer to $(X,\{u_{p,q}\}_{p,q\in P})$ as a product system in the sense of Fowler.
It can be directly verified that, if $X$ is a concrete product system in some ${\mathcal{B}}(H)$ such that $[X_{p}\cdot X_{q}]=X_{pq}$ for all $p,q\in P$, then the family $\{X_{p}\}_{p\in P}$ defines a product system in the sense of Fowler by considering the unitary maps $\{u_{p,q}\}_{p,q\in P}$ given by
@@
In this case the representations of $X$ are in bijection with the representations of $Y$. The reader may refer to [11, Section 2.3] for the full details.
@@ 3.1. Fock covariant representations
defined by the coaction by $G$ on ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$. A representation of $X$ that promotes to a representation of ${\mathcal{F}}{\mathcal{C}}X$ will be called a Fock covariant representation of $X$. We note that, the definition of the Fock covariant bundle does not depend on the choice of $G$. This follows by applying the first part of the proof of [52, Lemma 3.9] in this setting. Note also that since the bimodule properties are graded, every representation of ${\mathcal{F}}{\mathcal{C}}X$ is a representation of $X$.
@@
By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we have
@@
If $t$ is Fock covariant, then by definition $t_{\ast}$ factors through the canonical *-epimorphism $q_{\operatorname{c}}^{\operatorname{F}}\colon{\mathcal{T}}(X)\to{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$, and thus $\ker q_{\operatorname{c}}^{\operatorname{F}}\subseteq\ker t_{\ast}$. By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we also have $\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}$, and therefore
@@
For the converse, recall that the fixed point algebra $[{\mathcal{T}}(X)]_{e}$ is the inductive limit of $\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for $\cap$-closed families ${\mathcal{F}}\subseteq{\mathcal{J}}$. By the properties of the inductive limits we have
@@
Recall that we have $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ by Proposition 2.6. Therefore, if $r\notin\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then we have
@@
where in the first equality we used that we have $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. By injectivity of $\lambda_{r}$, we have $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. This concludes that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$, as required.
Now suppose that $t$ is Fock covariant. Recall that we have
@@
On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ then for every $\xi_{r}\in X_{r}$ we have
@@
where in the first equality we used that we have $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. Injectivity of $t$ implies that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. We conclude that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$ and thus
@@
For the converse, suppose that $t$ satisfies conditions (i) and (ii). By Proposition 3.1, it suffices to prove the inclusion $\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\subseteq\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for every finite $\cap$-closed ${\mathcal{F}}\subseteq{\mathcal{J}}$. If ${}{\mathcal{F}}=\{\emptyset\}$, then condition (i) for $\lambda$ and $t$ yields
@@
where we used the fact that the space $\mathbf{K}_{\emptyset,\hat{t}_{*}}=\mathbf{B}_{\{\emptyset\},\hat{t}_{\ast}}$ is a C*-subalgebra of $[{\mathcal{T}}(X)]_{e}$. Next suppose that =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ such that $\bigcup_{i=1}^{n}{\mathbf{x}}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$ such that
@@
For $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have
@@
where in the first equality we used Proposition 2.5, and on the second equality we used that we have $\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$ if $r\notin\mathbf{x}_{i}$ from Proposition 2.6. Condition (iii) now yields $\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$, and thus